1. Suppose f(K,L)=[L+K]3, what is the MRTS?

2. Suppose f(K,L)=K(1/2)+L(1/2) Find the marginal rate of technical substitution.

3. f(K,L)=K(1/2)+L(1/2) Find the marginal rate of technical substitution.
Answer #1

1- MPl = 3(L+K)^2

MPk= 3(L+K)^2

So,MRTS = MPl/MPk = 1

2. MPl= 1/2 L^ -1/2

MPk = 1/2 K^ -1/2

So,MRTS = MPl/MPk =( K/L )^1/2

2nd and 3rd question are same . Maybe you have done a typo while typing question.